# Think Smart, Not Hard: Difficulty Adaptive Reasoning for Large Audio Language Models

### v1

@@ 5 Related work
In recent years, advances in LLMs have driven the development of MLLMs, enhancing multimodal understanding. In the audio domain, LALMs such as Qwen2-Audio (Chu et al., 2024), Audio Flamingo (Kong et al., 2024), and SALMONN (Tang et al., 2024) handle basic understanding well but remain limited on complex tasks due to short outputs and lack of reasoning. Later models like Qwen2.5-Omni (Xu et al., 2025) and Kimi-Audio (KimiTeam et al., 2025) demonstrate some initial reasoning ability but still rely heavily on SFT, leaving outputs fixed and dependent on pretraining data, thus performing poorly in complex scenarios. To address this, some studies extend SFT, such as Audio-Reasoner (Xie et al., 2025), which uses large-scale CoT-annotated pairs to achieve reasoning via SFT. Others adopt RL, as in R1-AQA (Li et al., 2025) and Omni-R1 (Rouditchenko et al., 2025), which avoids using data with CoT, and promotes self-driven reasoning.
@@
Reasoning efficiency thus becomes a major challenge. Prior studies (Qu et al., 2025) highlight inefficiencies such as redundant content and overthinking. RL-based methods explore reward function designs to address this issue (Arora & Zanette, 2025; Aggarwal & Welleck, 2025; Shen et al., 2025). Most methods rely on setting a fixed length threshold for optimization. On the one hand, such approaches do not link reasoning length with problem type and therefore cannot adapt to all kinds of problems. On the other hand, the fixed threshold cannot change with the model’s evolving ability during training. Liu et al. (2025) summarizes these efforts and proposes a difficulty-aware dynamic approach that improves both performance and efficiency. However, this method depends on an additional independent dataset during training to measure the model’s capability, which not only increases computational cost but also makes the results sensitive to how this dataset is selected. In LALMs, Audio-Thinker (Wu et al., 2025) instead tackles the problem through a “when to think” mechanism, dividing tasks into those requiring reasoning and those that do not, but it does not further optimize the samples that require reasoning. In this work, we propose two difficulty-adaptive length-based rewards to enable efficient and effective reasoning in LALMs. Our method maintains comparable or even superior performance while reducing overall reasoning length, encouraging concise reasoning for simple questions, deeper reasoning for difficult ones, and progressive optimization as the model’s capability improves.
@@ Appendix A Appendix
A.1 Setup
@@
For the base models, we primarily adopt Qwen2-Audio-7B-Instruct and Qwen2.5-Omni-7B, which are among the most widely used open-source LALMs. These two models are also frequently used in prior work, facilitating fair and direct comparisons. For the training framework, we leverage the ms-swift (Zhao et al., 2025) and perform training on three A100-40G GPUs, where two GPUs are used for model training and one GPU is reserved for vLLM-based inference (Kwon et al., 2023). More training details and hyper-parameters can be found in the Appendix  A.4.
@@
For the training datasets, we use two in total: FS and AVQA (Yang et al., 2022). The latter is a subset of the former, while FS is constructed by augmenting AVQA with four additional datasets. For AVQA, we only keep the audio–text pairs and replace the word “video” in the questions with “audio.” For evaluation, we mainly test on MMAU-test-mini (Sakshi et al., 2025), with AirBench Foundation-Sound (Yang et al., 2024) as a secondary reference.
@@
GRPO has been extensively applied in both LLMs and MLLMs, achieving notable progress, and our implementation largely follows prior studies (DeepSeek-AI et al., 2025; Li et al., 2025). Compared with other RL methods, the key feature of GRPO is that it evaluates the policy model’s advantage using the average reward of in-group sampled outputs. Given an input question, a set of sampled responses for that question, and their rewards from the reward function, the advantage is calculated as follows:
$A_{i}=\frac{r_{i}-mean(\{r_{1},r_{2},\cdots,r_{G}\})}{std(\{r_{1},r_{2},\cdots,r_{G}\})}$ (8)
@@
To examine whether the effectiveness of CoT comes from explicit outputs or simply triggering implicit reasoning, we design two prompts, shown in Figure  3. For each sample, the model generates its answer within <answer> </answer>. In Prompt1, CoT is not required, while in Prompt2, it is generated within <think> </think>. A Format-Reward enforces this structure in Prompt2, whereas in Prompt1 it only regulates the answer format.
@@
A.2 Model-perspective Difficulty On MMAU
In Sakshi et al. (2025), question difficulty was annotated manually, where multiple experts assigned difficulty scores to each question. These labels are of very high quality, but the cost of such annotation is prohibitively high, and the labels are fixed, making it difficult to align them with the model’s evolving state across different training steps.
Therefore, in this subsection, we introduce model-perspective difficulty labels. Since each model has different capabilities, we aim to reflect an average level across models. To this end, we adopt four different models: Qwen2-Audio-7B-Instruct, Qwen2.5-Omni-7B, Kimi-Audio-7B-Instruct [^4] https://huggingface.co/moonshotai/Kimi-Audio-7B-Instruct, and Gemini2.5-Pro-0506. These models perform inference on MMAU-test-mini under the same random seed, and difficulty labels from the model’s perspective are assigned based on the number of models answering each question correctly. The detailed distribution of model-perspective difficulty labels is shown in Table  5.
@@ A.2 Model-perspective Difficulty On MMAU
A.3 Impact of Threshold Ratio on Our Proposed Reward
@@
Table 6: The performance of our proposed GRDR, GA2DR, and their variants with added length threshold ratios. Here, unmarked models correspond to $l_{min}=0$, and $\dagger$ denotes the model-perspective difficulty annotations.
@@
Table  6 reports the effects of adding a threshold ratio to our two proposed rewards. Models without special notation correspond to the case of $l_{min}=0$, which means no threshold ratio is applied. For the cases with $l_{min}=0.1$, this value is determined based on the average reasoning length of direct GRPO models. From a performance perspective, the overall results decline across both difficulty standards after introducing the threshold ratio. The only improvement appears on easy questions under model-perspective annotations, while all other cases show decreases. We attribute this to two main reasons. First, for some easy questions, the model already answers correctly at the start, receiving the maximum reward and thus maintaining its behavior. However, these samples still require optimization, and as a result, many easy questions remain unnecessarily long, with redundant reasoning or even error propagation from excessive reasoning. Second, for some hard questions, the model initially answers incorrectly without deeper reasoning, receiving the maximum penalty. Since nearby length ratios also correspond to maximum penalties, changes in length neither increase nor decrease the penalty, causing the model to stagnate without exploring deeper reasoning. In addition, when the model occasionally solves a hard question with a short but incorrect reasoning path, the large reward it receives may misguide the optimization process. Overall, instead of driving models toward excessively short and ineffective reasoning, the reward function (with maximum reward only at near-zero lengths) actually encourages models to learn what key reasoning steps to retain. Adding a threshold ratio, however, only introduces negative effects.
A.4 Hyper-parameters
@@
In this part, we provide detailed explanations of the hyperparameter settings used in the main experiments of Section  4, including the two proposed rewards and the basic TR hyperparameter settings.
@@
A.5 Example Analysis
@@
In this section, we present output examples from six different models on a sarcasm cause detection question in Figure  5. This question is labeled as Medium in the original MMAU-test-mini but is relabeled as Hard under our new scheme. The models include an advanced proprietary model, Gemini2.5-Pro, the prior work Audio-Reasoner, and the four main models used in our experiments. In the figure, the green text corresponds to the correct option for this example, the blue parts in each model’s reasoning indicate logical and necessary steps or key points for solving the question, and the red parts indicate repetition or redundancy, representing the parts that need optimization.

### v2

@@ 5 Related work
In recent years, advances in LLMs have driven the development of MLLMs, enhancing multimodal understanding. In the audio domain, LALMs such as Qwen2-Audio (DBLP:journals/corr/abs-2407-10759), Audio Flamingo (DBLP:conf/icml/KongGBPVC24), and SALMONN (DBLP:conf/iclr/TangYSC [^0] M024) handle basic understanding well but remain limited on complex tasks due to short outputs and lack of reasoning. Later models like Qwen2.5-Omni (DBLP:journals/corr/abs-2503-20215) and Kimi-Audio (DBLP:journals/corr/abs-2504-18425) demonstrate some initial reasoning ability but still rely heavily on SFT, leaving outputs fixed and dependent on pretraining data, thus performing poorly in complex scenarios. To address this, some studies extend SFT, such as Audio-Reasoner (DBLP:journals/corr/abs-2503-02318), which uses large-scale CoT-annotated pairs to achieve reasoning via SFT. Others adopt RL, as in R1-AQA (DBLP:journals/corr/abs-2503-11197) and Omni-R1 (DBLP:journals/corr/abs-2505-09439), which avoids using data with CoT, and promotes self-driven reasoning.
@@
Reasoning efficiency thus becomes a major challenge. Prior studies (DBLP:journals/corr/abs-2503-21614) highlight inefficiencies such as redundant content and overthinking. RL-based methods explore reward function designs to address this issue (DBLP:journals/corr/abs-2502-04463; DBLP:journals/corr/abs-2503-04697; DBLP:journals/corr/abs-2503-04472). Most methods rely on setting a fixed length threshold for optimization. On the one hand, such approaches do not link reasoning length with problem type and therefore cannot adapt to all kinds of problems. On the other hand, the fixed threshold cannot change with the model’s evolving ability during training. DBLP:journals/corr/abs-2505-15612 summarizes these efforts and proposes a difficulty-aware dynamic approach that improves both performance and efficiency. However, this method depends on an additional independent dataset during training to measure the model’s capability, which not only increases computational cost but also makes the results sensitive to how this dataset is selected. In LALMs, Audio-Thinker (wu2025audio) instead tackles the problem through a “when to think” mechanism, dividing tasks into those requiring reasoning and those that do not, but it does not further optimize the samples that require reasoning. In this work, we propose two difficulty-adaptive length-based rewards to enable efficient and effective reasoning in LALMs. Our method maintains comparable or even superior performance while reducing overall reasoning length, encouraging concise reasoning for simple questions, deeper reasoning for difficult ones, and progressive optimization as the model’s capability improves.
6 Limitations and Future work
In this work, we optimize reasoning across difficulty levels and achieve notable gains. But, limitations still remain. Our analysis of outputs from all models—including prior work and proprietary models—indicates that a strong CoT response should extract key information, perform structured reasoning, and deliver a clear final answer. Detailed analyses and examples are provided in Appendix  A.9 and Appendix  A.10. We also believe that incorporating rewards from external LLM APIs may further enhance the coherence and reliability of CoT outputs while preserving the intended reasoning paradigm, which we will explore in future work.
@@ Appendix A Appendix
A.1 Task Definition
Figure 4: An audio QA example from “FriendsFriends”. The top-left shows the question and options (green indicates the correct one), the right side presents the audio dialogue, and the bottom-left shows the output of our proposed method on Qwen2.5-Omni-7B.
A.2 Setup
@@
For the base models, we primarily adopt Qwen2-Audio-7B-Instruct and Qwen2.5-Omni-7B, which are among the most widely used open-source LALMs. These two models are also frequently used in prior work, facilitating fair and direct comparisons. For the training framework, we leverage the ms-swift (DBLP:conf/aaai/ZhaoHHWMZJWAWZC25) and perform training on three A100-40G GPUs, where two GPUs are used for model training and one GPU is reserved for vLLM-based inference (DBLP:conf/sosp/KwonLZ0ZY0ZS23). More training details and hyper-parameters can be found in the Appendix  A.8.
@@
For the training datasets, we use two in total: FS and AVQA (DBLP:conf/mm/Yang0DCHJ022). The latter is a subset of the former, while FS is constructed by augmenting AVQA with four additional datasets. For AVQA, we only keep the audio–text pairs and replace the word “video” in the questions with “audio.” For evaluation, we mainly test on MMAU-test-mini (DBLP:conf/iclr/SakshiTKSSNDGM25), with AirBench Foundation-Sound (DBLP:conf/acl/YangXLC0ZLLZZZ24) as a secondary reference.
@@
GRPO has been extensively applied in both LLMs and MLLMs, achieving notable progress, and our implementation largely follows prior studies (DBLP:journals/corr/abs-2501-12948; DBLP:journals/corr/abs-2503-11197). Compared with other RL methods, the key feature of GRPO is that it evaluates the policy model’s advantage using the average reward of in-group sampled outputs. Given an input question, a set of sampled responses for that question, and their rewards from the reward function, the advantage is calculated as follows:
$A_{i}=\frac{r_{i}-mean(\{r_{1},r_{2},\cdots,r_{G}\})}{std(\{r_{1},r_{2},\cdots,r_{G}\})}.$ (6)
@@
To examine whether the effectiveness of CoT comes from explicit outputs or simply triggering implicit reasoning, we design two prompts, shown in Figure  5. For each sample, the model generates its answer within <answer> </answer>. In Prompt1, CoT is not required, while in Prompt2, it is generated within <think> </think>. A Format-Reward enforces this structure in Prompt2, whereas in Prompt1 it only regulates the answer format.
@@
A.3 Model-perspective Difficulty On MMAU
In DBLP:conf/iclr/SakshiTKSSNDGM25, question difficulty was annotated manually, where multiple experts assigned difficulty scores to each question. These labels are of very high quality, but the cost of such annotation is prohibitively high, and the labels are fixed, making it difficult to align them with the model’s evolving state across different training steps.
Therefore, in this subsection, we introduce model-perspective difficulty labels. Since each model has different capabilities, we aim to reflect an average level across models. To this end, we adopt four different models: Qwen2-Audio-7B-Instruct, Qwen2.5-Omni-7B, Kimi-Audio-7B-Instruct [^4] https://huggingface.co/moonshotai/Kimi-Audio-7B-Instruct, and Gemini2.5-Pro-0506. These models perform inference on MMAU-test-mini under the same random seed, and difficulty labels from the model’s perspective are assigned based on the number of models answering each question correctly. The detailed distribution of model-perspective difficulty labels is shown in Table  9.
@@ A.3 Model-perspective Difficulty On MMAU
A.4 Impact of Threshold Ratio on Our Proposed Reward
@@
Table 10: The performance of our proposed GRDR, GA2DR, and their variants with added length threshold ratios on MMAU-Test-Mini. Here, unmarked models correspond to $l_{min}=0$, and $\dagger$ denotes the model-perspective difficulty annotations.
@@
Table 11: The performance of our proposed rewards, and their variants with added length threshold ratios on MMAU-v0515. Here, unmarked models correspond to $l_{min}=0$.
Models MMAU-Test-Mini-v0515
Sound Music Speech Avg
Based On Qwen2.5-Omni-7B
GRDR 83.48 70.35 75.97 76.60
+ $l_{min}=0.1$ 83.18 70.95 76.27 76.80
GA2DR 83.18 71.55 75.67 76.80
+ $l_{min}=0.1$ 81.08 72.15 77.77 77.00
Table 12: The performance of our proposed rewards, and their variants with added length threshold ratios on MMAR. Here, unmarked models correspond to $l_{min}=0$.
Models MMAR
Sound Music Speech Avg
Based On Qwen2.5-Omni-7B
GRDR 61.21 51.94 65.30 61.20
+ $l_{min}=0.1$ 63.63 52.91 65.98 63.00
GA2DR 64.84 54.85 65.30 62.90
+ $l_{min}=0.1$ 64.84 53.39 63.60 63.00
Table  10, Table  12, and Table  12 report the effects of adding a threshold ratio to our two proposed rewards on three benchmarks. Models without special notation correspond to the case of $l_{min}=0$, which means that no threshold ratio is applied. For the cases with $l_{min}=0.1$, this value is determined based on the average reasoning length of direct GRPO models.
From an overall performance perspective, models with a length threshold ratio achieve higher average scores on MMAU-Test-Mini-v0515 and MMAR, but show the opposite trend on MMAU-Test-Mini. Considering that the quality of the Q&As and audio files in MMAU-Test-Mini is relatively poor, the performance on the latter two benchmarks is therefore more convincing. The fact that introducing a length threshold leads to better results on most benchmarks further indicates that our length-based reward curve tends to infinitely optimize toward a completion length of zero, which may cause the CoT content of certain samples to become less effective, ultimately resulting in a decrease in overall performance.
From the perspective of different difficulty levels, this part mainly focuses on the results from the two versions of MMAU-Test-Mini. We observe that the model with the added threshold performs better on easy and hard questions, while the model without the threshold achieves better results on medium questions. This outcome is closely related to our threshold setting: the threshold value was determined based on the average reasoning length of the direct GRPO model on MMAU-Test-Mini. Since more than half of the questions in the dataset are labeled as medium difficulty, the threshold essentially aligns with this group. As a result, medium questions— which could have benefited from further length optimization—received less optimization, leading to relatively lower performance. In contrast, for easy and hard questions, introducing the threshold brings two benefits: for easy questions, it prevents excessively short reasoning that may reduce CoT effectiveness; for hard questions, it increases the upper bound of reasoning length, enabling deeper exploration of problem-solving strategies. Consequently, the model achieves better performance on both easy and hard questions.
In addition, when focusing on GRDR and GA2DR, we can observe that the latter achieves better performance on two benchmarks. This is because GA2DR defines difficulty levels at the batch level, which aligns with the batch-wise optimization process during model backpropagation, making this approach more effective. Moreover, its attention mechanism jointly considers both the textual question and the characteristics of the audio itself.
Furthermore, previous studies (peng2025simko) have shown that models trained with reinforcement learning tend to exhibit improvements in pass@1 but declines in pass@k, largely due to the instability of rollout samples. Our GA2DR method effectively mitigates this issue, since all rollout samples within a group share the same question and audio input, thereby avoiding inconsistencies caused by rollout variance.
In summary, combining GA2DR with a length threshold provides the most balanced and effective optimization strategy. The specific threshold value, however, should be determined based on the characteristics of each individual task.
A.5 Detailed length statistics of different models on MMAU
In this section, we provide the raw reasoning-length data for all models used in our main experiments on MMAU-Test-Mini, without any post-processing. All lengths are measured using the Qwen2-Audio tokenizer. Note that Qwen2-Audio and Qwen2.5-Omni share the same tokenizer, so using the former does not introduce any error when evaluating the latter. For Gemini2.5, however, we cannot obtain its tokenizer because it is a closed-source model. Although using the Qwen2-Audio tokenizer may introduce some deviation, the reasoning lengths of Gemini2.5 are far greater than those of all our methods, so this small discrepancy does not affect the clearly observable relative differences.
Table 13: The output length statistics of models on MMAU-test-mini with different base models and reward settings, reported under both the human-perspective difficulty annotations and the assigned model-perspective difficulty annotations. Here, $\dagger$ denotes the model-perspective difficulty annotations.
Models MMAU-Test-Mini
Easy $\text{Easy}^{\dagger}$ Medium $\text{Medium}^{\dagger}$ Hard $\text{Hard}^{\dagger}$
Baseline Models
Qwen2-Audio-7B-Instruct 59.79 65.48 67.66 65.26 70.18 68.50
Qwen2.5-Omni-7B 151.80 89.76 80.19 121.90 121.57 134.67
Kimi-Audio-7B-Instruct 94.99 21.79 29.69 59.03 25.01 82.38
Gemini2.5-Pro-0506 759.60 525.85 542.17 663.58 628.32 768.81
Based On Qwen2-Audio-7B-Instruct
GRPO 41.76 42.26 42.55 43.29 43.89 42.97
+ TR 42.11 43.55 43.77 44.42 46.32 44.31
+ GRDR 27.76 30.41 31.59 31.49 32.26 30.88
Based On Qwen2.5-Omni-7B
GRPO 185.38 108.08 99.74 153.90 146.18 164.92
+ TR 183.65 110.20 103.10 153.49 145.76 165.46
+ GRDR 54.49 51.69 41.70 56.05 68.32 56.73
  xxx + $l_{min}=0.1$ 124.62 89.10 85.66 110.00 107.54 116.96
+ GA2DR 54.58 52.84 48.57 53.92 66.48 57.58
  xxx + $l_{min}=0.1$ 115.38 88.80 85.34 105.89 105.91 109.68
A.6 Reasoning length vs. Accuracy
In this section, we analyze the relationship between output length and accuracy for the GRDR and GA2DR methods on MMAU-Test-Mini. From Figure  7 to Figure  10, samples are grouped into length intervals of similar size, and correctly answered samples within each interval are further divided by difficulty level, indicated by increasingly darker shades of blue.
Figure 7: The trend of length and accuracy for the GRDR with human-perspective difficulty.
Figure 8: The trend of length and accuracy for the GRDR with model-perspective difficulty.
Figure 9: The trend of length and accuracy for the GA2DR with human-perspective difficulty.
Figure 10: The trend of length and accuracy for the GA2DR with model-perspective difficulty.
A.7 Training Curves
In this section, we present the training curves of GRDR and GA2DR on Qwen2.5-Omni, including grad norm, KL divergence, and reward trends in Figure  11 and Figure  12. All curves are directly extracted from the TensorBoard logs recorded during training, with a smoothing factor of 0.8 to clearly reveal their variations and overall trajectories.
Overall, both methods converge stably without exhibiting extreme values in gradients or KL divergence, and no numerical spikes are observed. This indicates that our training procedures are stable and that the proposed methods are well-behaved in practice.
((a)) Grad Norm
((b)) KL Divergence
((c)) Reward
Figure 11: Training curves of gradient norm, KL divergence, and reward on GRDR.
((a)) Grad Norm
((b)) KL Divergence
((c)) Reward
Figure 12: Training curves of gradient norm, KL divergence, and reward on GA2DR.
A.8 Hyper-parameters
@@
In this part, we provide detailed explanations of the hyperparameter settings in Table  14, including the two proposed rewards and the basic TR hyperparameter settings.
@@
A.9 Case Study: Qualitative Analysis of Output Paradigms across Models
In this section, we analyze output paradigms across models using 20 MMAU-test-mini questions on Dissonant Emotion Interpretation, focusing on sarcasm cause detection. The corresponding audios often involve multiple speakers and complex environments, making the task challenging and well-suited for qualitative analysis of different model outputs. In this section, in addition to the main models discussed above, we also include a Cold-Start GRPO model based on Qwen2-Audio-7B-Instruct, with detailed settings provided in the Appendix  A.2.
The detailed results are shown in Table  15, which include ACC and reasoning length on the subset mentioned above. In the second part, “Based on Qwen2-Audio-7B-Instruct,” we compare our three GRPO models with Audio-Reasoner. GRPO and GRPO with GRDR match Audio-Reasoner in performance while producing much shorter outputs, whereas the GRPO model after Cold-Start produces reasoning lengths close to Audio-Reasoner but performs worst. We believe this is mainly because the base model is weaker and can only imitate the surface structure of advanced model paradigms, showing that learning paradigms through Cold-Start is not effective in all cases and can even yield negative effects. In the “Based on Qwen2.5-Omni-7B” part of the table, performance differences are minor, mainly because the Qwen2.5-Omni base model already has some reasoning ability and shows good performance on complex questions. However, with our two proposed rewards, the models maintain their original performance while significantly shortening reasoning length, greatly improving reasoning efficiency.
Model ACC Avg-Length
Advanced Proprietary Model
Gemini2.5-Pro-0506 95 931.8
Based On Qwen2-Audio-7B-Instruct
Audio-Reasoner 75 547.1
GRPO 75 50.7
  $x$ + GRDR 70 38.5
  $x$ + Cold-Start SFT 55 541.4
Based On Qwen2.5-Omni-7B
GRPO 100 109.5
  $x$ + GRDR 95 56.2
  $x$ + GRDR ($l_{min}=0.1$) 90 94.1
  $x$ + GA2DR 100 56.9
  $x$ + GA2DR ($l_{min}=0.1$) 95 94.2
In addition to analyzing performance and reasoning length above, we also compared the outputs of different models on these 20 questions in detail. From the results, we found that models with better performance usually follow a complete reasoning process, which includes first grounding the audio and providing a corresponding caption—equivalent to identifying the known conditions in the problem—then reasoning based on these conditions, analyzing each option step by step, and finally giving the answer. This approach helps LALMs make better use of known information together with their broad pretrained knowledge to perform reasonable reasoning while also improving readability. However, weaker models, even when using this structure, often make mistakes in the first step of extracting known information. These errors propagate through the reasoning process, leading to wrong answers and lower performance. For this part, we also provide an example of different models’ outputs in Appendix  A.10 as reference.
A.10 Example Analysis
In this section, we present output examples from six different models on a sarcasm cause detection task from Figure  13 to Figure  15. This question is labeled as Medium in the original MMAU-test-mini but is relabeled as Hard under our new scheme. The models include an advanced proprietary model, Gemini2.5-Pro, the prior work Audio-Reasoner, and the four main models used in our experiments. In the figure, the green text corresponds to the correct option for this example, the blue parts in each model’s reasoning indicate logical and necessary steps or key points for solving the question, and the red parts indicate repetition or redundancy, representing the parts that need optimization.